solve the system using the elimination method with work shown.
2x+6y=17
2x-10y=9

Answer by mananth(16949)

2 x + 6 y = 17 .............1
2 x -10 y = 9 .............2
Eliminate y
multiply (1)by 5
Multiply (2) by 3
10 x 30 y = 85
6 x -30 y = 27
Add the two equations
16 x = 112
/ 16
x = 7
plug value of x in (1)
2 x + 6 y = 17
14 + 6 y = 17
6 y = 17-14
6 y = 3
y = 0.50